# A Modified 3D Biconical Outflow Model: Spatial Constraints on AGN-driven Outflows

### v1

@@ 2.3 Seeing Effect
where $F_{c}(x,y),V_{c}(x,y)$, and $\sigma_{c}(x,y)$ are convolved flux, velocity, and velocity dispersion, respectively. In the bottom row of Figure 2, we show an example of convolved 2D maps. The red horizontal bar on the upper right represents the applied seeing size in this example (FWHM =0.8=0.8″). The seeing convolution blurs the central bicone structure, so the final images are more like a circular shape in all 2D maps. We will discuss the seeing effect in Section 4.3 and 5.3, focusing on implications for the outflow size measurements in IFU data.
2.4 Mock Data Compatibility
We update our model to calculate 2D maps in an (x, y) coordinate system defined in arcsec units based on the input redshift and physical size of the outflow, while Bae16 model adopted relative scales normalized by the outflow size. This approach offers several advantages: (1) the outflow model uses the same pixel units as the disk model, making it convenient to combine the two models; (2) mock data can be directly compared with real observational data. Accordingly, we specify the mock data by pixel size and field-of-view (FoV) parameters, and all calculations are performed using this pixel scale from given redshift and physical length parameters ($R_{out}$, $r_{1/2}$, and $r_{t}$).
Figure 2: 2D map examples of the modified model. From left to right, three 2D maps in each row present flux, velocity, and velocity dispersion of the corresponding model. First row: Bae16 biconical model. This model has the same geometry as Figure 1. Second row: Disk model from GalPak3D. Third row: The bicone + disk combined model. Last row: The combined model after seeing convolution. Red dashed circles at the center and red bars on the top right show the FWHM seeing size. We also plot 3″-aperture size as gray circles, which is the fiber size of SDSS spectra.
3 Mock Data Analysis
In this Section, we describe our main analysis procedures based on mock data generated from the modified outflow model. First, we generate mock SDSS data extracted from the central 3″-aperture, and compare them with the observed [O III] kinematics of a large sample of SDSS type 2 AGNs in Section 3.1. In Section 3.2, we produce mock IFU data to analyze outflow sizes and 2D distributions of flux, velocity, and velocity dispersion.
3.1 Mock SDSS Data
Bae16 integrated the [O III] velocity offset and dispersion within the outflow radius and compared them with the observed VVD distribution of $\sim$39,000 SDSS type 2 AGNs from Woo16. Here, we use our modified model to reexamine the comparison with the observed VVD distribution.
@@ 3.1 Mock SDSS Data
3.1.1 Model Grids for the Observed VVD
@@
3.1.2 Mock VVD from Monte Carlo Simulation
@@
We first take PDFs of various parameters, such as redshift, stellar mass, disk-to-cone flux ratio, and extinction, from the observed SDSS type 2 AGNs (see Figure 3). The PDF of redshift (0.04 to 0.28 with 0.04 steps) is directly applied to the MC simulations. The observed disk-to-cone flux ratio is determined by the flux ratio between broad/wing and narrow/core components when the [O III] line is fitted with a double Gaussian. If the [O III] line is fitted with a single Gaussian, we classify the emission as either disk-only or outflow-only based on the difference between stellar and [O III] velocity dispersion. Specifically, when $|\sigma_{\mathrm{[OIII]}}-\sigma_{*}|$ is less than 30 km s-1, we classify this single Gaussian as disk-only emission, i.e., $f_{disk}/f_{cone}=\infty$, otherwise as outflow-only emission, i.e., $f_{disk}/f_{cone}=0$. We also estimate the PDF of the dust extinction factor by using observed $\alpha/H$$\beta$ with $R^{A}_{V}=4.5$ (Vogt13). Note that we restrict the dust extinction rate to smaller than 80% (or $A>0.2$) since a higher extinction rate results in too broad mock VVD distributions along the V[OIII] axis, which are inconsistent with the observed distribution. Furthermore, an extinction rate higher than 80% is an extreme case, corresponding to $A_{\lambda 5007}\gtrsim$ 1.75 mag.
@@
The stellar mass is taken from the MPA-JHU catalog  [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/ and divided into five bins ($\rm{log}\;M_{*}=$9.7 to 11.3 with 0.4 step). We calculate the four remaining disk parameters, i.e., $r_{1/2}$ (Dutton11), $V_{max,rot}$ (Torres-Flores11), $\sigma_{0}$ (Oh22), and $r_{t}$ (Amorisco10), using empirical relations with stellar mass. Thus, once the stellar mass is selected among the five values, these four disk parameters are determined accordingly. Note that we restrict our sample to type 2 AGNs in the MC simulations due to the large uncertainties in stellar mass for type 1 AGNs.
@@
For the outflow model, we first adopt random inclinations and fixed PAs (PAcone=0 and PAdisk=90) for both the bicone and dust plane. The inclination of the dust plane is given from 7.5 to 172.5 with 15 step, and its PDF is proportional to the solid angle of the normal vector integrated from 0 to 360 of PA. Since we only deal with type 2 AGNs, the bicone inclination is restricted to less than 40, i.e., 7.5, 22.5, and 37.5, and its PDF is also proportional to the solid angle of the bicone axis. Then we assume that the PDF for five outflow size values (0.3, 1, 3, 6, 9 kpc) follows a Gaussian distribution with a mean of zero and a standard deviation of 2 kpc. Note that the outflow size has little effect on the kinematics in mock SDSS data.
@@
Our main goal is to constrain two parameters, i.e., the launching velocity and opening angle of the outflow model. For these two parameters, we assume that their PDFs follow a truncated Gaussian function. We first set the range of the launching velocity to $100,\;\infty)$ km s-1  and restrict the outer opening angle values from 30 to 70 with a 10 step. We then run multiple MC simulations by changing the mean and standard deviation of the Gaussian to find the best $V_{max}$ and $\theta_{out}$ distributions that generate the best-matching mock VVD with the observed VVD. Each MC simulation generates 100,000 pairs of velocity offset and dispersion.
3.1.3 Kullback-Leibler Divergence and $D_{max}$
@@
To quantitatively compare the mock VVDs with the observed VVD, we apply two statistical quantities that assess proximity between two 2D distributions. We first calculate Kullback-Leibler divergence ($D_{KL}(P||Q)$, KLDIV) as follows:
@@
where $P(v_{obs},\sigma_{obs})$ and $Q(v_{int},\sigma_{int})$ are probability distributions of the observed and mock VVD, respectively. In brief, the Kullback-Leibler divergence is a measure of a statistical distance between two PDFs derived from information theory, i.e., the smaller the Kullback-Leibler divergence is, the more similar the two PDFs are.
@@
Note that we add random error smoothing to all mock VVDs based on typical SDSS uncertainties of velocity offset and dispersion to account for the observational uncertainties. To remove variations induced by this error smoothing, we repeat this process 100 times in each grid of $V_{max}$ and $\theta_{out}$ PDFs and take the mean and standard deviation of Kullback-Leibler divergence and $D_{max}$ values. If the mean Kullback-Leibler divergence (or $D_{max}$) for given parameter distributions is notably smaller than for the others, by more than its standard deviation, then that mock VVD distribution is considered a better reproduction of the observed data.
Figure 3: Histograms of physical properties of Woo16 type 2 AGNs. From left to right, each histogram shows redshift, stellar mass, flux ratio between disk and outflow components, and extinction factor based on Balmer decrement.
3.2 Mock IFU Data
@@
In our previous IFU studies, we developed a kinematic approach to measure the outflow size and applied it to AGNs at $z<0.3$ (Karouzos16a; Kang18; Bae17; Luo21; Kim23). The outflow region is defined as where the [O III] velocity dispersion is more enhanced than stellar velocity dispersion, which traces the gravitational potential of the host galaxies. Specifically, we first plot a radial profile of ratios between the [O III] and stellar velocity dispersion ($\sigma_{\mathrm{[OIII]}}/\sigma_{*}$) and define the outflow size as a radius where the ratio becomes unity. Then we quadratically subtract the seeing size ($\sigma_{g}$) from the outflow size to correct the seeing effect.
We apply the same procedure to mock IFU data and compare the measured outflow size with the input outflow size. Since we do not have stellar velocity dispersion in our model, we alternatively adopt the [O III] velocity dispersion extracted within a 3″-aperture of the convolved disk-only maps. This velocity dispersion of the disk model mimics the [O III] narrow/core component, which is an optimal proxy of stellar velocity dispersion (e.g., Greene05).
Another widely used method for outflow size measurements is W80-based method (e.g., Harrison14; Wylezalek20; Ruschel-Dutra21), where W80 is defined as a [O III] line width including 80% of the total line flux. Similar to the $\sigma_{\mathrm{[OIII]}}$-based method, the W80-based method defines the outflow region where the W80 is broader than a certain criterion, frequently set to 600 km s-1. Thus, the outflow size is determined as the radius of the furthest pixel from the center with W80 larger than 600 km s-1.
@@
4 Result
4.1 Location of Model Grids in VVD Diagram
@@
Figure 4: Model grids from the modified outflow model. Each color indicates a different launching velocity, 500 km s-1(cyan), 1000 km s-1(yellow), and 2000 km s-1(red): (a) The observed [O III] VVD distribution of type 2 AGNs (Woo16). This color map is also plotted as contours and gray color maps in the other three panels; (b) Model grids with (solid lines) and without (dashed lines) the disk. Each line shows the change in kinematics depending on extinction at a fixed inclination; (c) Model grids showing the effects of bicone inclinations and extinction rates when the disk component is included. Solid lines indicate variations with extinction at a fixed bicone inclination, while dotted lines indicate variations with inclination at a fixed extinction rate; (d) Model grids with different opening angles and extinction rates. Each solid and dotted line indicates the same bicone opening angle and extinction rate, respectively.
@@
We investigate the dependency of outflow parameters on the location in the VVD diagram by comparing model grids with observations. The observed VVD distribution of local type 2 AGNs in Figure 4(a) is characterized by a V-shaped envelope and more AGNs with blueshifted [O III] than redshifted (Woo16).
@@
First, we examine the effect of the rotating disk emission by comparing model grids with/without the disk component (Figure 4(b)), finding that both the integrated velocity offset and dispersion ($v_{int}$ and $\sigma_{int}$) decrease by including a disk component. This is because the disk component usually has a smaller velocity offset and dispersion than the outflow, resulting in smaller velocity offset and dispersion values in the combined model than in the outflow-only model. Thus, the modified model with a disk component requires a higher launching velocity ($\lesssim$ 1500 km s-1) to reproduce a similar VVD range, while most of AGNs in the VVD diagram were explained by $V_{max}\lesssim 1000$ km s-1(orange dashed lines) in the original model (Bae16).
@@
Second, we revisit the effect of the bicone inclination and the dust extinction (Figure 4(c)). We find that both $\sigma_{int}$ and $v_{int}$ increase at a more inclined bicone, which is consistent with Bae16. For higher extinction, $\sigma_{int}$ decreases while $v_{int}$ increases because of the less contribution of the receding cone. Notably, the angle of the V-shape becomes smaller than that of the outflow-only model since the ratio of $|v_{int}|/\sigma_{int}$ decreases due to the zero velocity offset of the disk component.
@@
Lastly, we investigate the effect of the opening angle (Figure 4(d)). We find that a wider opening angle generates a larger $\sigma_{int}$ due to a broader range of LOS velocities, while $v_{int}$ is almost constant regardless of the opening angle. Overall, the effect of the opening angle is consistent with Bae16.
4.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
We constrain the distributions of $V_{max}$ and $\theta_{out}$ parameters by comparing mock VVD distributions from the MC simulations with the observed VVD distribution. We conduct 315 realizations of MC simulations using Gaussian distributions, for which the mean and standard deviation of $V_{max}$ change from 150 to 450 km s-1and from 250 to 450 km s-1, respectively, in 50 km s-1steps. Mean and standard deviation of the $\theta_{out}$ parameter are set to 30, 40, 50 and 5, 10, 20, respectively. Note that the adopted mean and standard deviation are different from those of the resulting PDFs, as we are using a truncated Gaussian distribution.

### v2

@@ II.3 Seeing Effect
where $F_{c}(x,y),V_{c}(x,y)$, and $\sigma_{c}(x,y)$ are convolved flux, velocity, and velocity dispersion, respectively. In the bottom row of Figure 2, we show an example of convolved 2D maps. The red horizontal bar on the upper right represents the applied seeing size in this example (FWHM =0.8=0.8″). The seeing convolution blurs the central bicone structure, so the final images are more like a circular shape in all 2D maps. We will discuss the seeing effect in Section IV.3 and V.3, focusing on implications for the outflow size measurements in IFU data.
II.4 Mock Data Compatibility
We update our model to calculate 2D maps in an (x, y) coordinate system defined in arcsec units based on the input redshift and physical size of the outflow, while the Bae and Woo (2016) model adopted relative scales normalized by the outflow size. This approach offers several advantages: (1) the outflow model uses the same pixel units as the disk model, making it convenient to combine the two models; (2) mock data can be directly compared with real observational data. Accordingly, we specify the mock data by pixel size and field-of-view (FoV) parameters, and all calculations are performed using this pixel scale from given redshift and physical length parameters ($R_{out}$, $r_{1/2}$, and $r_{t}$).
Figure 2: Examples of 2D maps produced by the modified model. From left to right, each column shows the flux, velocity, and velocity dispersion maps. The first row presents the biconical outflow model of Bae and Woo (2016). The second row shows the rotating disk model generated with GalPak3D. The third row displays the combined bicone + disk model, and the last row presents the combined model after applying the seeing convolution. Red dashed circles at the center, and red bars in the upper-right corner indicate the FWHM of the seeing. Gray circles denote the 3″ aperture corresponding to the SDSS fiber size.
III Mock Data Analysis
In this Section, we describe our main analysis procedures based on mock data generated from the modified outflow model. First, we generate mock SDSS data extracted from the central 3″-aperture, and compare them with the observed [O III] kinematics of a large sample of SDSS type 2 AGNs in Section III.1. In Section III.2, we produce mock IFU data to analyze outflow sizes and 2D distributions of flux, velocity, and velocity dispersion.
III.1 Mock SDSS Data
Bae and Woo (2016) integrated the [O III] velocity offset and dispersion within the outflow radius and compared them with the observed VVD distribution of $\sim$39,000 SDSS type 2 AGNs from Woo et al. (2016). Here, we use our modified model to reexamine the comparison with the observed VVD distribution.
@@ III.1 Mock SDSS Data
III.1.1 Model Grids for the Observed VVD
@@
III.1.2 Mock VVD from Monte Carlo Simulation
@@
We first take PDFs of various parameters, such as redshift, stellar mass, disk-to-cone flux ratio, and extinction, from the observed SDSS type 2 AGNs (see Figure 3). The PDF of redshift (0.04 to 0.28 with 0.04 steps) is directly applied to the MC simulations. The observed disk-to-cone flux ratio is determined by the flux ratio between broad/wing and narrow/core components when the [O III] line is fitted with a double Gaussian. If the [O III] line is fitted with a single Gaussian, we classify the emission as either disk-only or outflow-only based on the difference between stellar and [O III] velocity dispersion. Specifically, when $|\sigma_{\mathrm{[OIII]}}-\sigma_{*}|$ is less than 30 km s-1, we classify this single Gaussian as disk-only emission, i.e., $f_{disk}/f_{cone}=\infty$, otherwise as outflow-only emission, i.e., $f_{disk}/f_{cone}=0$. We also estimate the PDF of the dust extinction factor by using observed $\alpha/H$$\beta$ with $R^{A}_{V}=4.5$ (Vogt et al., 2013). Note that we restrict the dust extinction rate to smaller than 80% (or $A>0.2$) since a higher extinction rate results in too broad mock VVD distributions along the V[OIII] axis, which are inconsistent with the observed distribution. Furthermore, an extinction rate higher than 80% is an extreme case, corresponding to $A_{\lambda 5007}\gtrsim$ 1.75 mag.
@@
The stellar mass is taken from the MPA-JHU catalog  [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/ and divided into five bins ($\rm{log}\;M_{*}=$9.7 to 11.3 with 0.4 step). We calculate the four remaining disk parameters, i.e., $r_{1/2}$ (Dutton et al., 2011), $V_{max,rot}$ (Torres-Flores et al., 2011), $\sigma_{0}$ (Oh et al., 2022), and $r_{t}$ (Amorisco and Bertin, 2010), using empirical relations with stellar mass. Thus, once the stellar mass is selected among the five values, these four disk parameters are determined accordingly. Note that we restrict our sample to type 2 AGNs in the MC simulations due to the large uncertainties in stellar mass for type 1 AGNs.
@@
For the outflow model, we first adopt random inclinations and fixed PAs (PAcone=0 and PAdisk=90) for both the bicone and dust plane. The inclination of the dust plane is given from 7.5 to 172.5 with a 15 step, and its PDF is proportional to the solid angle of the normal vector integrated from 0 to 360 of PA. Since we focus on type 2 AGNs, we restrict the bicone inclination to less than 40 (7.5, 22.5, and 37.5), and its PDF is also proportional to the solid angle of the bicone axis. Models with higher bicone inclinations, i.e., more aligned bicone axis with the LOS, are likely to be observed as type 1 AGNs, as their accretion disk and broad line region (BLR) are readily exposed to the observer (e.g., Marin, 2014). Then we assume that the PDF for five outflow size values (0.3, 1, 3, 6, 9 kpc) follows a Gaussian distribution with a mean of zero and a standard deviation of 2 kpc. Note that the outflow size has little effect on the kinematics in mock SDSS data.
@@
Our main goal is to constrain two parameters, i.e., the launching velocity and opening angle of the outflow model. For these two parameters, we assume that their PDFs follow a truncated Gaussian function. We first set the range of the launching velocity to $100,\;\infty)$ km s-1 and restrict the outer opening angle values from 30 to 70 with a 10 step based on Fischer et al. (2013). Note that larger outer opening angles are excluded due to their physically extreme geometry, while smaller opening angles are excluded since they do not improve the main results. We then run multiple MC simulations by changing the mean and standard deviation of the Gaussian to find the best $V_{max}$ and $\theta_{out}$ distributions that generate the best-matching mock VVD with the observed VVD. Each MC simulation generates 100,000 pairs of velocity offset and dispersion.
III.1.3 Kullback-Leibler Divergence and $D_{max}$
@@
To quantitatively compare the mock VVDs with the observed VVD, we apply two statistical quantities that assess proximity between two 2D distributions. We first calculate Kullback-Leibler (KL) divergence ($D_{KL}(P||Q)$, Kullback and Leibler, 1951) as follows:
@@
where $P(v_{obs},\sigma_{obs})$ and $Q(v_{int},\sigma_{int})$ are probability distributions of the observed and mock VVD, respectively. In brief, the KL divergence is a measure of a statistical distance between two PDFs derived from information theory, i.e., the smaller the KL divergence is, the more similar the two PDFs are.
@@
Note that we add random error smoothing to all mock VVDs based on typical SDSS uncertainties of velocity offset and dispersion to account for the observational uncertainties. To remove variations induced by this error smoothing, we repeat this process 100 times in each grid of $V_{max}$ and $\theta_{out}$ PDFs and take the mean and standard deviation of KL divergence and $D_{max}$ values. If the mean KL divergence (or $D_{max}$) for given parameter distributions is notably smaller than for the others, by more than its standard deviation, then that mock VVD distribution is considered a better reproduction of the observed data.
Figure 3: Histograms of physical properties of the type 2 AGN sample from Woo et al. (2016). From left to right, the panels show the distributions of redshift, stellar mass, flux ratio between disk and outflow components, and extinction factor derived from the Balmer decrement.
III.2 Mock IFU Data
@@
In our previous IFU studies, we developed a kinematic approach to measure the outflow size and applied it to AGNs at $z<0.3$ (Karouzos et al., 2016; Kang and Woo, 2018; Bae et al., 2017; Luo et al., 2021; Kim et al., 2023). The outflow region is defined as where the [O III] velocity dispersion is more enhanced than stellar velocity dispersion, which traces the gravitational potential of the host galaxies. Specifically, we first plot a radial profile of ratios between the [O III] and stellar velocity dispersion ($\sigma_{\mathrm{[OIII]}}/\sigma_{*}$) and define the outflow size as a radius where the ratio becomes unity. Then we quadratically subtract the seeing size ($\sigma_{g}$) from the outflow size to correct the seeing effect.
We apply the same procedure to mock IFU data and compare the measured outflow size with the input outflow size. Since we do not have stellar velocity dispersion in our model, we alternatively adopt the [O III] velocity dispersion extracted within a 3″-aperture of the convolved disk-only maps. This velocity dispersion of the disk model mimics the [O III] narrow/core component, which is an optimal proxy of stellar velocity dispersion (e.g., Greene and Ho, 2005).
Another widely used method for outflow size measurements is W80-based method (e.g., Harrison et al., 2014; Wylezalek et al., 2020; Ruschel-Dutra et al., 2021), where W80 is defined as a [O III] line width including 80% of the total line flux. Similar to the $\sigma_{\mathrm{[OIII]}}$-based method, the W80-based method defines the outflow region where the W80 is broader than a certain criterion, frequently set to 600 km s-1. Thus, the outflow size is determined as the radius of the furthest pixel from the center with W80 larger than 600 km s-1.
@@
IV Result
IV.1 Location of Model Grids in VVD Diagram
@@
Figure 4: Model grids with a fixed launching velocity of 500 km s-1 (cyan), 1000 km s-1 (yellow), and 2000 km s-1 (red). (a) Observed [O III] VVD distribution of 39,000 type 2 AGNs from Woo et al. (2016); the color scale indicates the number of AGNs in each bin. This distribution is also shown as contours and gray heat maps in panels (b)–(d). (b) Model grids with (solid lines) and without (dashed lines) the disk component. Each line traces the change in kinematics as a function of the extinction at a fixed inclination. (c) Model grids showing the effects of bicone inclination and extinction when the disk component is included. Solid lines indicate variations with extinction at a fixed inclination, while dotted lines indicate variations with inclination at a fixed extinction. (d) Model grids with different opening angles and extinction values. For clarity, solid and dotted lines denote models with fixed opening angles and fixed extinction values, respectively.
@@
We investigate the dependency of outflow parameters on the location in the VVD diagram by comparing model grids with observations. The observed VVD distribution of local type 2 AGNs (Figure 4(a)) is characterized by a V-shaped envelope and a larger number of blueshifted than redshifted [O III] sources (Woo et al., 2016).
@@
First, we examine the influence of the rotating disk component by comparing model grids with and without disk emission (Figure 4(b)). Including the disk component reduces both the integrated velocity offset and dispersion ($v_{\rm int}$ and $\sigma_{\rm int}$), because the disk typically exhibits smaller velocity offsets and dispersions than the outflow. As a result, the combined model requires a higher launching velocity up to 1500 km s-1 to reproduce the observed VVD distribution, whereas the original outflow-only model (Bae and Woo, 2016) could explain most AGNs with $V_{\rm max}$ up to 1000 km s-1 (orange dashed lines).
@@
Second, we revisit the effects of bicone inclination and dust extinction (Figure 4(c)). Both $v_{\rm int}$ and $\sigma_{\rm int}$ increase with larger inclination angles, consistent with Bae and Woo (2016). Increasing extinction decreases $\sigma_{\rm int}$ but increases $v_{\rm int}$, owing to the reduced contribution from the receding cone. Notably, the $|v_{\rm int}|/\sigma_{\rm int}$ ratio is generally smaller than in the outflow-only model because the disk component has zero intrinsic velocity offset.
@@
Lastly, we investigate the effect of the opening angle (Figure 4(d)). A wider opening angle produces a larger $\sigma_{\rm int}$ due to the broader distribution of LOS velocities, while $v_{\rm int}$ remains nearly constant. Overall, the dependence on opening angle is consistent with the trends reported by Bae and Woo (2016).
IV.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
We constrain the distributions of $V_{\rm max}$ and $\theta_{\rm out}$ by comparing the simulated VVD distributions with the observed VVD distribution. We perform multiple MC realizations, adopting Gaussian PDFs in which the mean and standard deviation of $V_{\rm max}$ range from 150 to 450 km s-1 and from 250 to 450 km s-1, respectively, in steps of 50 km s-1. For $\theta_{\rm out}$, we use mean values of $30^{\circ}$, $40^{\circ}$, and $50^{\circ}$, and standard deviations of $5^{\circ}$, $10^{\circ}$, and $20^{\circ}$. We note that the adopted means and standard deviations do not correspond exactly to those of the resulting PDFs because the distributions are truncated Gaussians. We then calculate 4D arrays containing 315 values ($7\times 5\times 3\times 3$) for the KL divergence and $D_{max}$, and identify the best PDFs where these values are minimized. As a result, KL divergence is minimized when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(250 km s-1, 400 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 10). $D_{max}$ exhibits a minimum value when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(350 km s-1, 300 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 5).